A credit union client deposits $4,000 in an account earning 2.5% interest, compounded daily. What will the balance of the account be at the end of 21
years?
Enter the answer in dollars and cents, and round to the nearest cent, if needed. Do not include the dollar sign. For example, if the answer is $0.61, only the
number 0.61 should be entered.
Balance $ ____________ after 21 years.

Answer by LinnW(1048)

$4000 for 21 years compounded daily at 2.5% annual rate.
The number of compounding periods is 365*21 = 7665
The interest rate for a compounding period is 2.5% / 365
= 0.025/365
So the answer is 4000*(1 + 0.025/365)^7665 = 6761.71